Unit 29 Homework Assignment  Physics141-Summer 2009

Unit 29 homework after session 1

• Read chapter 35 in the textbook Understanding Physics.

• Work textbook problems 35-21 and 35-26. Unit 29 homework after session 2

• Read chapter 35 in the text book Understanding Physics.

• Work textbook problems 35-22, 35-25, and 35-29.

All Unit 29 homework problems and the unit 29 activity guide are due on the first day of Unit 30.